Let (x1,x2,x3) and (y1,y2,y3) represent the coordinates with respect to the bases B1 = {(1, 0, 0),(1, 1, 0),(0, 0, 1)},B2 = {(1, 0, 0),(0, 1, 2),(0, 2, 1)}. If Q(x) = x1^2-4x1x2+2x2x3+x2^2+x3^2 , find the representation of Q in terms of (y1,y2,y3)
1
Expert's answer
2020-07-12T17:47:53-0400

Let "e_{1} = eS_1, e_2 = eS_2, e_2 = e_1S"", e = {(1;0;0), (0;1;0), (0;0;1)}" and "A = \begin{pmatrix}
 1& -2 & 0\\
 -2& 1 & 1\\
 0 & 1 & 1
\end{pmatrix}" is the mathix of Q(x).

Then "S_1 = \begin{pmatrix}
 1& 1 & 0\\
 0& 1 & 0\\
 0 & 0 & 1
\end{pmatrix}" , "S_2 = \begin{pmatrix}
 1& 0 & 0\\
 0& 1 & 2\\
 0 & 2 & 1
\end{pmatrix}" , "e = e_1S_1^{-1}" , "e_2 = eS_2 = e_1S_1^{-1}S_2"


That is "S = S_1^{-1}S_2"


"S_1^{-1} = \begin{pmatrix}
 1& -1 & 0\\
 0& 1 & 0\\
 0 & 0 & 1
\end{pmatrix}" "S =S_1^{-1}S_2 = \begin{pmatrix}
 1& -1 & 0\\
 0& 1 & 0\\
 0 & 0 & 1
\end{pmatrix} \cdot \begin{pmatrix}
 1& 0 & 0\\
 0& 1 & 2\\
 0 & 2 & 1
\end{pmatrix}
= \begin{pmatrix}
 1& -1 & -2\\
 0& 1 & 2\\
 0 & 2 & 1
\end{pmatrix}"


"A' = S^TAS = \begin{pmatrix}
 1& 0 & 0\\
 -1& 1 & 2\\
 -2 & 2 & 1 
\end{pmatrix} \cdot \begin{pmatrix}
 1& -2 & 0\\
 -2& 1 & 1\\
 0 & 1 & 1
\end{pmatrix} \cdot \begin{pmatrix}
 1& -1 & -2\\
 0& 1 & 2\\
 0 & 2 & 1
\end{pmatrix}" "="


"= \begin{pmatrix}
 1& -2 & 0\\
 -3& 5 & 3\\
 -6 & 7 & 3
\end{pmatrix} \cdot \begin{pmatrix}
 1& -1 & -2\\
 0& 1 & 2\\
 0 & 2 & 1
\end{pmatrix}" "= \begin{pmatrix}
 1& -3 & -6\\
 -3& 14 & 19\\
 -6 & 19 & 29
\end{pmatrix}" .


So, "Q(y) = y_1^2 - 6y_1y_2 -12y_1y_3 + 14y_2^2 + 38y_2y_3 +29y_3^2"